# Double Robustness of Local Projections

### v1

@@ 6 Conclusion
We have shown that, while conventional LP inference is robust to surprisingly large amounts of misspecification, VAR inference is much more fragile. Conventional VAR confidence intervals suffer from severe coverage distortions even for misspecification that is small, difficult to rule out ex ante using economic theory, and difficult to detect ex post with VAR model specification tests. These analytical results demonstrate that—when the goal is to construct valid confidence intervals for impulse responses, as opposed to minimizing MSE—the smaller bias of LPs documented in simulations by Li, Plagborg-Møller, and Wolf (2024) is more valuable than the smaller variance enjoyed by VAR estimators.
@@
A practical take-away for LP analysis is that researchers should control for lags of the data that are strong predictors of the outcome and impulse variables. This is important even if the researcher directly observes a near-perfect proxy for the shock of interest. However, it is not necessary to get the lag length exactly right to achieve correct coverage, and conventional information criteria (such as AIC) suffice to select an appropriate lag length for LP in our framework. Our results complement the finding of Montiel Olea and Plagborg-Møller (2021) that lag-augmented LP confidence intervals are also more robust than VAR intervals to persistence in the data and to the length of the impulse response horizon.
@@
Is there a way forward for VAR inference? The conventional VAR confidence interval is valid if the estimation lag length is sufficiently large, but our results show that this can only happen by making the interval equivalent with the LP interval. Using a smaller lag length combined with a bias-aware critical value typically leads to wider confidence intervals than LP. Another option would be to estimate VARMA models rather than pure VARs, though this would be computationally expensive, and the bias-variance trade-off relative to LPs is unclear. In principle, VAR procedures may work better under additional restrictions on the misspecification, such as shape restrictions on the impulse response functions. [^13] Given any convex parameter space for the misspecification MA polynomial $\alpha(L)$, the worst-case bias of the VAR estimator (see Proposition 3.2) can be computed using convex programming. However, it appears that detailed application-specific restrictions would be required to generate a negligible worst-case bias, since we have shown that the least favorable misspecification in our baseline analysis is economically realistic. Rather than restricting the parameter space, an alternative would be to weaken the coverage requirement, e.g., only requiring a certain coverage probability on average over a set of horizons (Armstrong, Kolesár, and Plagborg-Møller, 2022), or by changing the target for inference from the true impulse response function to a smooth projection of this function (Genovese and Wasserman, 2008). Finally, a subjectivist Bayesian VAR modeler need only worry about our negative results if their prior on potential misspecification attaches significant weight to MA processes that imply large VAR biases.
@@ A.1 Least favorable misspecification
Figure A.1 plots some numerical examples of the least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha^{\dagger}_{\ell,h,M}L^{\ell}$ discussed in Section 4.3. We focus here on the univariate local-to-AR(1) model from Section 2, though unreported numerical experiments suggest that the qualitative features mentioned below also apply to multivariate models. Recall that the least favorable MA coefficients depend on the horizon $h$ of interest, while $M$ only influences the overall scale of the coefficients, and not their shape as a function of the lag $\ell$. The figure shows that the shape of the coefficients either takes the form of a hump or a single zig-zag pattern, with the largest absolute value of the coefficients generally occurring at $\ell=h$. Notice that we can flip the signs of all coefficients without changing the bias.
@@ A.2 More efficient bias-aware confidence interval
Generalizing the bias-aware VAR confidence interval in Section 4.4, consider a bias-aware confidence interval that is centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$ from 4.2:
$\operatorname{CI}_{B}(\hat{\theta}_{h}(\omega);M)\equiv\left[\hat{\theta}_{h}(\omega)\pm\operatorname{cv}_{1-a}\left(\frac{(1-\omega)M\lambda}{\sqrt{1+\omega^{2}\lambda^{2}}}\right)\sqrt{(1+\omega^{2}\lambda^{2})\operatorname{aVar}(\hat{\delta}_{h})/T}\right],$
where $\lambda\equiv\sqrt{\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})-1}$. This interval equals the conventional LP interval when $\omega=1$ and the bias-aware VAR interval when $\omega=0$.
@@
The result follows from Propositions 3.1, 3.2, 3.1 and 4.1 and the same calculations as in the proof of 4.2. ∎
If we choose the weight $\omega$ to minimize confidence interval length, the resulting bias-aware interval tends to be nearly as long as the LP interval. The length-optimal averaging weight $\omega=\omega^{*}$ is given by
$\omega^{*}\equiv\operatorname*{argmin}_{\omega\in[0,1]}\;\operatorname{cv}_{1-a}\left(\frac{(1-\omega)M\lambda}{\sqrt{1+\omega^{2}\lambda^{2}}}\right)\sqrt{1+\omega^{2}\lambda^{2}}.$
Figure A.3 shows this optimal weight as a function of $M$ and the relative asymptotic standard deviation of the VAR and LP estimators, while Figure A.3 shows the length of the resulting optimal bias-aware confidence interval relative to the length of the conventional LP interval. We see that, for $M\geq 2$, there is little gain from reporting the optimal bias-aware interval rather than the LP interval, regardless of the relative precision of VAR and LP. An additional observation is that, for $M\geq 2$, the length-optimal $\omega^{*}$ is numerically close to the MSE-optimal weight $M^{2}/(1+M^{2})$ derived in 4.2.
@@
Appendix B Proofs of propositions
@@
Define
$\operatorname{aBias}(\hat{\delta}_{h})\equiv\operatorname*{trace}\left\{S^{-1}\Psi_{h}H\sum_{\ell=1}^{\infty}\alpha_{\ell}DH^{\prime}(A^{\prime})^{\ell-1}\right\}-e_{i^{*},n}^{\prime}\sum_{\ell=1}^{h}A^{h-\ell}H\alpha_{\ell}e_{j^{*},m}.$
@@
By Propositions 3.1 and 3.2 and 3.1,
$\hat{\beta}_{h}-\hat{\delta}_{h}$ $=-T^{-\zeta}\operatorname{aBias}(\hat{\delta}_{h})+R_{T}+o_{p}(T^{-\zeta}+T^{-1/2}),$
@@
[…]
Consider the following three cases:
  1. 1.
    $\zeta\in(1/4,1/2)$: In this case
    $\lim_{T\to\infty}P\left(\frac{\sqrt{T}|\hat{\beta}_{h}-\hat{\delta}_{h}|}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}}>z_{1-a/2}\right)$ $=\lim_{T\to\infty}P\left(\frac{T^{\zeta}|\hat{\beta}_{h}-\hat{\delta}_{h}|}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}}>\frac{z_{1-a/2}}{T^{1/2-\zeta}}\right)$
    $=P\left(\frac{|\operatorname{aBias}(\hat{\delta}_{h})|}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}}>0\right)$
    $=1.$
    where the last inequality uses the fact that $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$.
  2. 2.
    $\zeta>1/2$: In this case
[…]
    and thus
    $\lim_{T\to\infty}P\left(\frac{\sqrt{T}|\hat{\beta}_{h}-\hat{\delta}_{h}|}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}}>z_{1-a/2}\right)=a.$
  3. 3.
    $\zeta=1/2$: In this case
[…]
    Thus,
    $\lim_{T\to\infty}P\left(\frac{\sqrt{T}|\hat{\beta}_{h}-\hat{\delta}_{h}|}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}}>z_{1-a/2}\right)=r\left(\frac{b_{h}}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})-1}};z_{1-a/2}\right),$
    where $b_{h}\equiv\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$ and $r(b;c)\equiv P_{Z\sim N(0,1)}(|Z+b|>c)=\Phi(-c-b)+\Phi(-c+b)$, as in 3.2.
In summary, the three cases above show that
$\lim_{T\to\infty}P\left(\frac{\sqrt{T}|\hat{\beta}_{h}-\hat{\delta}_{h}|}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})}}>z_{1-a/2}\right)=\lim_{T\to\infty}r\left(\frac{T^{1/2-\zeta}b_{h}}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})-1}};z_{1-a/2}\right).\qed$
B.4 Proof of Proposition 4.1
Rewrite the numerator of $b_{h}$ as
@@
$\operatorname*{trace}\left\{S^{-1}\Psi_{h}H\sum_{\ell=1}^{\infty}\alpha_{\ell}DH^{\prime}(A^{\prime})^{\ell-1}\right\}-e_{i^{*},n}^{\prime}\sum_{\ell=1}^{h}A^{h-\ell}H\alpha_{\ell}e_{j^{*},m}=\sum_{\ell=1}^{\infty}\operatorname*{trace}\left\{\Xi_{h,\ell}D^{-1/2}\alpha_{\ell}D^{1/2}\right\},$
with
$\Xi_{h,\ell}\equiv D^{1/2}H^{\prime}(A^{\prime})^{\ell-1}S^{-1}\Psi_{h}HD^{1/2}-\mathbbm{1}(\ell\leq h)D^{-1/2}e_{j^{*},m}e_{i^{*},n}^{\prime}A^{h-\ell}HD^{1/2}.$
By Cauchy-Schwarz, we can bound the squared value of the numerator of $b_{h}$ as follows:
$\left(\sum_{\ell=1}^{\infty}\operatorname*{trace}\left\{\Xi_{h,\ell}D^{-1/2}\alpha_{\ell}D^{1/2}\right\}\right)^{2}$ $\leq\left(\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{h,\ell}^{\prime}\Xi_{h,\ell})\right)$
×$\qquad\times\underbrace{\left(\sum_{\ell=1}^{\infty}\operatorname*{trace}\{(D^{-1/2}\alpha_{\ell}D^{1/2})^{\prime}D^{-1/2}\alpha_{\ell}D^{1/2}\}\right)}_{=\|\alpha(L)\|^{2}}.$
Moreover, the bound is achieved by choosing $\alpha_{\ell}\propto\Xi_{h,\ell}^{\prime}$ for $\ell=1,2,\dots$. Hence, the statement of the proposition follows if we can show that

### v2

@@ 6 Conclusion
Our theoretical results suggest the following practical take-aways:
  1. 1.
    When the goal is to construct confidence intervals for impulse responses that have accurate coverage in a wide range of empirically relevant DGPs—as opposed to minimizing MSE—then the smaller bias of LPs documented in simulations by Li, Plagborg-Møller, and Wolf (2024) is more valuable than the smaller variance enjoyed by VAR estimators.
  2. 2.
@@
    Researchers who use LP should control for those lags of the data that are strong predictors of the outcome or impulse variables. This is important even if the researcher directly observes a near-perfect proxy for the shock of interest. However, it is not necessary to get the lag length exactly right to achieve correct coverage. To select the number of lags to control for in the LP, it suffices to run a VAR in all variables used in the analysis and select the lag length that minimizes conventional information criteria (such as AIC). Our results complement the finding of Montiel Olea and Plagborg-Møller (2021) that lag-augmented LP confidence intervals are also more robust than VAR intervals to persistence in the data and to the length of the impulse response horizon.
  3. 3.
    There is no free lunch for VARs: if an estimated VAR yields confidence intervals that are substantially narrower than the corresponding LP intervals, we recommend increasing the VAR lag length until that is no longer the case, to guarantee robust confidence interval coverage. Conventional tests of correct VAR specification do not suffice to guard against coverage distortions.
@@
Is there a way forward for VAR inference, beyond just including a large number of lags? We showed how to construct a VAR confidence interval with a bias-aware critical value that robustly controls coverage, but found that it will typically lead to wider confidence intervals than LP. Another option would be to estimate VARMA models rather than pure VARs, though this would be computationally expensive, and the bias-variance trade-off relative to LPs is unclear. In principle, VAR procedures may work better under additional restrictions on the misspecification, such as shape restrictions on the impulse response functions. [^12] Given any convex parameter space for the misspecification MA polynomial $\alpha(L)$, the worst-case bias of the VAR estimator (see Proposition 3.2) can be computed using convex programming. However, it appears that detailed application-specific restrictions would be required to generate a negligible worst-case bias, since we have shown that the least favorable misspecification in our baseline analysis is economically plausible. Rather than restricting the parameter space, future research could instead investigate weakening the coverage requirement, e.g., only requiring a certain coverage probability on average over a set of horizons (Armstrong, Kolesár, and Plagborg-Møller, 2022), or by changing the target for inference from the true impulse response function to a smooth projection of this function (Genovese and Wasserman, 2008). Finally, a subjectivist Bayesian VAR modeler need only worry about our negative results if their prior on potential misspecification attaches significant weight to MA processes that imply large VAR biases.
@@ A.1 Least favorable misspecification
Figure A.1 plots some numerical examples of the least favorable MA polynomial $\alpha^{\dagger}(L;h,M)=\sum_{\ell=1}^{\infty}\alpha^{\dagger}_{\ell,h,M}L^{\ell}$ discussed in Section 4.2. We focus here on a univariate local-to-AR(1) model $y_{t}=\rho y_{t-1}+[1+T^{-1/2}\alpha(L)]\varepsilon_{t}$, though unreported numerical experiments suggest that the qualitative features mentioned below also apply to multivariate models. Recall that the least favorable MA coefficients depend on the horizon $h$ of interest, while $M$ only influences the overall scale of the coefficients, and not their shape as a function of $\ell$. The figure shows that the shape of the coefficients either takes the form of a hump or of a single zig-zag pattern, with the largest absolute value of the coefficients generally occurring at $\ell=h$. Notice that we can flip the signs of all coefficients without changing the bias.
@@ A.2 More efficient bias-aware confidence interval
Generalizing the bias-aware VAR confidence interval in Section 4.3, consider a bias-aware confidence interval that is centered at the model averaging estimator $\hat{\theta}_{h}(\omega)=\omega\hat{\beta}_{h}+(1-\omega)\hat{\delta}_{h}$ from Corollary 4.2:
$\operatorname{CI}_{B}(\hat{\theta}_{h}(\omega);M)\equiv\left[\hat{\theta}_{h}(\omega)\pm\operatorname{cv}_{1-a}\left(\frac{(1-\omega)M\tau}{\sqrt{1+\omega^{2}\tau^{2}}}\right)\sqrt{(1+\omega^{2}\tau^{2})\operatorname{aVar}(\hat{\delta}_{h})/T}\right],$
where $\tau\equiv\sqrt{\operatorname{aVar}(\hat{\beta}_{h})/\operatorname{aVar}(\hat{\delta}_{h})-1}$. This interval equals the conventional LP interval when $\omega=1$ and the bias-aware VAR interval when $\omega=0$.
@@
The result follows from Propositions 3.1, 3.2, A.2 and 4.1 and the same calculations as in the proof of Corollary 4.2. ∎
Even if we choose the weight $\omega$ to minimize confidence interval length, the resulting bias-aware interval tends to be nearly as long as the LP interval. The length-optimal weight $\omega=\omega^{*}$ is given by
$\omega^{*}\equiv\operatorname*{argmin}_{\omega\in[0,1]}\;\operatorname{cv}_{1-a}\left(\frac{(1-\omega)M\tau}{\sqrt{1+\omega^{2}\tau^{2}}}\right)\sqrt{1+\omega^{2}\tau^{2}}.$
Figure A.3 shows this optimal weight as a function of $M$ and the relative asymptotic standard deviation of the VAR and LP estimators, while Figure A.3 shows the length of the resulting optimal bias-aware confidence interval relative to the length of the conventional LP interval. We see that, for $M\geq 2$, there is little gain from reporting the optimal bias-aware interval rather than the LP interval, regardless of the relative precision of VAR and LP. An additional observation is that, for $M\geq 1.5$, the length-optimal $\omega^{*}$ is numerically close to the MSE-optimal weight $M^{2}/(1+M^{2})$ derived in Corollary 4.2.
@@
A.3 Covariance structure of LP and VAR estimators
The following result provides the asymptotic variance-covariance matrix of the LP and VAR estimators in the general multi-dimensional set-up of Section 4.4. Define $\Psi_{i^{*},j^{*},h}$ as in Proposition 3.2, but making the dependence on $(i^{*},j^{*})$ explicit in the notation.
Corollary A.2.
Impose Assumption 2.1, with part (iii) holding for all shock indices $j_{1}^{*},\dots,j_{k}^{*}$. Then for any $a,b\in\{1,\dots,k\}$,
$\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\beta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ $=\mathbbm{1}(j_{a}^{*}=j_{b}^{*})\sigma_{j_{a}^{*}}^{-2}\left(\psi_{a,b}+\sum_{\ell=1}^{\min\{h_{a},h_{b}\}}e_{i_{a}^{*},n}^{\prime}A^{h_{a}-\ell}\Sigma(A^{\prime})^{h_{b}-\ell}e_{i_{b}^{*},n}\right),$
$\operatorname{aCov}(\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ $=\mathbbm{1}(j_{a}^{*}=j_{b}^{*})\sigma_{j_{a}^{*}}^{-2}\psi_{a,b}+\operatorname*{trace}\left(\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}\Sigma\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}^{\prime}S^{-1}\right),$
$\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})$ $=\operatorname{aCov}(\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}}),$
where
$\psi_{a,b}\equiv e_{i_{a}^{*},n}^{\prime}A^{h_{a}}\overline{H}_{j_{a}^{*}}\overline{D}_{j_{a}^{*}}\overline{H}_{j_{a}^{*}}^{\prime}(A^{\prime})^{h_{b}}e_{i_{b}^{*},n},$
and the “$\operatorname{aCov}$” notation is understood to refer to elements of the asymptotic variance-covariance matrix in (4.2). In particular, $\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}}-\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}})=0$.
Proof.
See Section B.9. ∎
Appendix B Proofs
Lemmas whose name begins with “E” can be found in the Supplemental Appendix.
@@ Appendix B Proofs
Define $\tilde{\alpha}_{\ell}=D^{-1/2}\alpha_{\ell}D^{1/2}$ for all $\ell\geq 1$. Notice that $\|\alpha(L)\|^{2}=\sum_{\ell=1}^{\infty}\|\tilde{\alpha}_{\ell}\|^{2}$.
By Proposition 3.2, we have $\operatorname{aBias}(\hat{\delta}_{i^{*},j^{*},h})=\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{i^{*},j^{*},h,\ell}\tilde{\alpha}_{\ell})$, where
$\Xi_{i^{*},j^{*},h,\ell}\equiv D^{1/2}H^{\prime}(A^{\prime})^{\ell-1}S^{-1}\Psi_{i^{*},j^{*},h}HD^{1/2}-\mathbbm{1}(\ell\leq h)D^{-1/2}e_{j^{*},m}e_{i^{*},n}^{\prime}A^{h-\ell}HD^{1/2}.$
Since $\operatorname*{trace}(\Xi_{i^{*},j^{*},h,\ell}\tilde{\alpha}_{\ell})=\operatorname*{vec}(\Xi_{i^{*},j^{*},h,\ell})^{\prime}\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime})$, we can write
$\operatorname{aBias}(\hat{\boldsymbol{\delta}})=\sum_{\ell=1}^{\infty}\Upsilon_{\ell}\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime}),$
@@
$\Upsilon_{\ell}\equiv\Big{(}\operatorname*{vec}(\Xi_{i_{1}^{*},j_{1}^{*},h_{1},\ell}),\dots,\operatorname*{vec}(\Xi_{i_{k}^{*},j_{k}^{*},h_{k},\ell})\Big{)}^{\prime}\in\mathbb{R}^{k\times m^{2}}.$
Hence,
@@
$\max_{\alpha(L)\colon\|\alpha(L)\|\leq M}\|R\operatorname{aBias}(\hat{\boldsymbol{\delta}})\|^{2}=\max_{\{\tilde{\alpha}_{\ell}\}_{\ell=1}^{\infty}\colon\sum_{\ell=1}^{\infty}\|\tilde{\alpha}_{\ell}^{\prime}\|^{2}\leq M^{2}}\left\|\sum_{\ell=1}^{\infty}R\Upsilon_{\ell}\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime})\right\|^{2}.$
LABEL:thm:bias_general_aux1 shows that the final expression above equals $M^{2}\lambda_{\max}(\sum_{\ell=1}^{\infty}R\Upsilon_{\ell}\Upsilon_{\ell}^{\prime}R^{\prime})$ (the lemma only explicitly considers the case $M=1$, but the general case then follows from the homogeneity of degree 1 of the norm). Finally, Lemma B.1 below shows that $\sum_{\ell=1}^{\infty}\Upsilon_{\ell}\Upsilon_{\ell}^{\prime}=\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})$. This completes the proof of the proposition. The proof of LABEL:thm:bias_general_aux1 shows that the maximum above is achieved when $\operatorname*{vec}(\tilde{\alpha}_{\ell}^{\prime})\propto\Upsilon_{\ell}^{\prime}v$ (with the constant of proportionality being independent of $\ell$ and chosen to satisfy the norm constraint), where $v$ is the eigenvector corresponding to the largest eigenvalue of $R[\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}})]R^{\prime}$. In the univariate case $k=1$, this reduces to expression (4.1) in Section 4.2. ∎
Lemma B.1.
Under the assumptions of Proposition 4.2, and using the notation in the proof of that proposition, we have